There are 18 animals in the barnyard. Some are chickens and some are cows. I counted 50 legs in all. How many of the animals are chickens and how many are cows?
Answer by Nate(3500)

There are 18 animals in the barnyard. Some are chickens and some are cows. I counted 50 legs in all. How many of the animals are chickens and how many are cows?
Define variables:
c=number of chickens
o=number of cows
Define equations:
c+o=18
2c+4o=50
Combine:
-2c-2o=-36
+
2c+4o=50
=
2o=14
o=7
Substitute:
c+o=18
c+7-18
c=11
Eleven chickens and seven cows are in barnyard!